OpenStudy (anonymous):

Find the equation of the line that passes through (3,4) and m = -4.

OpenStudy (zzr0ck3r):

y = mx+b we know m y = -4x+b now plug in the pointto solve for b 4 = -4*3 = b can you do th rest?

OpenStudy (anonymous):

That's a good way to do it^. You can also put what you have into Point-Slope form. Then change it into slope-intercept.

OpenStudy (zzr0ck3r):

that should say 4 = -4*3+b

OpenStudy (anonymous):

y - y1 = m(x - x1) y - 4 = -4(x - 3) And just solve for y.

OpenStudy (anonymous):

y=16-4x

OpenStudy (anonymous):

y=y=-4x+16

OpenStudy (anonymous):

Yep, nice work(: